# From User Surveys to Telemetry-Driven Agents: Exploring the Potential of Personalized Productivity Solutions

### v1

@@ 7. Limitations
Despite the valuable insights gained from this study, some limitations should be considered when interpreting the findings and generalizing the results. One key limitation was the sample size and diversity, which may not have been large enough or sufficiently diverse in terms of demographics, professional backgrounds, and technology expertise to capture the full range of preferences, experiences, and productivity patterns among users. This lack of diversity could limit the generalizability of our findings to broader populations and contexts. The study’s duration may also not have been long enough to fully assess the effects of the personalized productivity agent, dashboard visualization, and narrative summary, as participants may require more time to adapt to these tools and strategies. It is essential to acknowledge the potential influence of using different prompts on the generated outputs by GPT-4. As the AI model relies heavily on the input prompts provided, variations in these prompts can lead to dramatically different results, which could affect the overall consistency and reliability of the findings. The language model’s performance is highly sensitive to the phrasing and structure of the prompts, and even subtle changes might result in significantly different responses. Consequently, the outcomes derived from multiple iterations of prompts may not always be directly comparable, and this limitation should be considered when interpreting the results and drawing conclusions based on the AI-generated content. Also of note are the occasional hallucinations exhibited by GPT-4 and its poor ability to perform calculations accurately. As a result, we refrained from asking the agent to generate additional summary information or perform complex calculations, which could have potentially provided more insights into the productivity patterns. In our survey methodology (Appendix B), detailed questions about the interactive agent were followed by queries on the solutions’ strengths, weaknesses, and suggestions (Q28, Q29). This sequence could potentially induce a response bias, where participants’ views on the agent’s performance might have disproportionately influenced their assessment of its strengths and weaknesses. While we aimed for a comprehensive understanding of user experiences and randomized other aspects of the study including Q27 which explicitly asks for the benefits and drawbacks of each presentation method immediately after interacting with that method, future studies may benefit from randomizing the order of these questions to mitigate any potential bias.
@@
The telemetry data extracted from Viva Insights provided valuable information on participant work patterns but did not capture a complete picture. Furthermore, there might have been external factors or variables, such as changes in workload, organizational structure, or personal circumstances, that may affect productivity and thus are essential to be considered. In light of these limitations, it is essential to approach the study’s findings with caution and acknowledge that the results may not be universally applicable. Future research could address these limitations by employing larger and more diverse samples, extending the study duration, incorporating additional measures of productivity, and controlling for potential confounding factors. By doing so, researchers can continue to advance our understanding of the factors that contribute to productivity enhancement and the role of personalized productivity agents in this process. Finally, it’s important to consider the extent to which the AI productivity agent’s design was influenced by the user preferences identified in Phase 1. While there is a possibility that a closer alignment might have affected the comparative results, our design was a conscientious effort to balance user needs with technical feasibility. Future research could explore a more in-depth alignment of these insights, potentially offering a refined understanding of how well the AI agent meets user expectations. This ongoing process highlights the iterative nature of developing user-centric AI tools.
@@ 8. Ethical Considerations
In this study, we adhered to ethical standards to protect the rights, privacy, and well-being of our participants. Informed consent was obtained from all participants, who were provided with a comprehensive overview of the study’s objectives, procedures, potential benefits, and risks. Anonymity and confidentiality were maintained by using unique identification codes for participant data, securely storing all information on encrypted servers, and limiting access to authorized research team members.
As the use of LLMs becomes increasingly prevalent in various aspects of our lives, it is essential to consider the ethical implications of integrating these models with other data sources, such as telemetry data. The combination of these technologies may lead to significant advancements in productivity and user experience. However, it also raises several ethical concerns, including data privacy, security, potential biases, surveillance, and the unintended consequences of automation. To address these concerns, we must maintain an ongoing conversation about the ethical implications of integrating large language models with other data. This includes addressing potential biases in AI models, ensuring transparency in data processing and decision-making, and establishing strong privacy and security protocols. It is also essential to involve a diverse range of stakeholders, such as ethicists, policymakers, and representatives from affected communities, in the development and deployment of these technologies. By fostering a proactive and responsible approach to the ethical challenges presented by the integration of large language models, we can harness the power of these technologies in a manner that respects individual rights and promotes the well-being of society as a whole.
@@ 9. Conclusion
This paper presents a comprehensive study focused on understanding user preferences for productivity agents and designing personalized solutions to cater to these preferences. By conducting a user survey and analyzing the findings, we were able to identify key features and behaviors that contribute to user satisfaction and comfort when interacting with a productivity agent. Subsequently, we developed an AI-based productivity agent and compared its performance with alternative modes of telemetry data presentation, such as dashboard visualization and personalized narratives. The results from this study provide valuable insights into the factors that enhance productivity and user experience, as well as areas for improvement in the design and development of personalized productivity agents. By incorporating the lessons learned from this study, future work can focus on refining and optimizing productivity-enhancing tools and solutions, ultimately leading to improved efficiency and user-centric experiences for information workers.
@@ Appendix C Prompts for Extracting Data
User Prompt for Page 1: Extract the organized meetings, invited meetings and meeting habits (along with %) and list them. If you cannot find these information, do not mention anything about it in the response. DO NOT LIST ANY NAMES OF INDIVIDUALS. List the information in this format: ‘Meetings organized in the past four weeks: X, Meetings invited to in the past four weeks: X, Meeting habits based on meetings you attended: - Advanced notice: X%, - High attendance: X%, - No overlap with other meetings: X%, - Ended on time: X%,- Joined on time: X% - During working hours: X% - Meeting was $\leq$ 1 hour: X% - RSVP’d to invite: X% - You didn’t multitask: X% - Added a teams link: X%’
@@ Appendix E Prompt for Personalized Narrative and Interactive Chat Agent
Appendix H User Survey
@@
Making progress or completing scheduled tasks
347 (95.6%)
Solving novel, unexpected tasks
278 (76.6%)
Helping a colleague accomplish a task
262 (72.2%)
Learning new techniques or information
250 (68.9%)
Staying up-to-date with messaging and emails
153 (42.1%)
Effectively managing and participating in meetings
149 (41.0%)
Establishing new connections or expanding my professional network
111 (30.6%)
Other
14 (3.9%)
@@
Yes, definitely
142 (39.1%)
Yes, somewhat
138 (38.0%)
Neutral
57 (15.7%)
No, not really
17 (4.7%)
No, not at all
9 (2.5%)
@@
Moderate impact
117 (41.8%)
Minimal impact
73 (26.1%)
Significant impact
54 (19.3%)
Unsure
27 (9.6%)
Other
9 (3.2%)
@@
Casual and friendly
46 (16.4%)
Formal and professional
23 (8.2%)
It depends on the context
95 (33.9%)
A balance of both
99 (35.4%)
No strong preference
17 (6.1%)
@@
Professional and respectful
71 (25.4%)
Direct and straightforward
69 (24.6%)
Adaptive
55 (19.6%)
Friendly and warm
50 (17.9%)
Engaging and motivational
19 (6.8%)
Unsure
12 (4.3%)
Other
4 (1.4%)
@@
Direct and to the point
99 (35.4%)
Chatty and conversational
12 (4.3%)
A balance of both
86 (30.7%)
It depends on the context
80 (28.6%)
No strong preference
3 (1.1%)
@@
Moderate Personalization
109 (38.9%)
High Personalization
100 (35.7%)
Minimal Personalization
34 (12.1%)
No personalization
15 (5.4%)
Unsure
22 (7.9%)
@@
Yes, completely comfortable
46 (16.4%)
Somewhat comfortable
106 (37.9%)
Neutral
61 (21.8%)
Uncomfortable
58 (20.7%)
Unsure
9 (3.2%)
@@
Yes, completely comfortable
121 (43.2%)
Somewhat comfortable
84 (30.0%)
Neutral
36 (12.9%)
Uncomfortable
31 (11.1%)
Unsure
8 (2.9%)
@@
Proactive suggestions
169 (60.4%)
Only specific requests
77 (27.5%)
Not at all
7 (2.5%)
Unsure
19 (6.8%)
Other
8 (2.9%)
@@
Knowledgeable
249 (88.9%)
Trustworthy
210 (75.0%)
Transparent
191 (68.2%)
Responsive
184 (65.7%)
Adaptable
172 (61.4%)
Professional
131 (46.8%)
Friendly
104 (37.1%)
Humor
92 (32.9%)
Empathetic
81 (28.9%)
No strong preference
4 (1.4%)
Other
13 (4.6%)
@@
Assistant
214 (76.4%)
Expert
157 (56.1%)
Coworker
96 (34.3%)
Mentor
93 (33.2%)
Friend
65 (23.2%)
Supervisor
13 (4.6%)
Unsure
7 (2.5%)
Other
2 (0.7%)

### v2

@@ 7. Limitations & Future Work
In addition, while we developed prompts through iterative testing, we did not conduct systematic prompt engineering to optimize the balance between accuracy and computational cost. Also of note are the occasional hallucinations exhibited by GPT-4 and its poor ability to perform calculations accurately. As a result, we refrained from asking the agent to generate additional summary information or perform complex calculations, which could have potentially provided more insights into the productivity patterns. While we did not systematically analyze the chat transcripts for hallucinations, it is noteworthy that none of the participants reported experiencing incorrect or hallucinated information in their open-ended survey responses about the agent’s performance. In our survey methodology (Appendix B), detailed questions about the interactive agent were followed by queries on the solutions’ strengths, weaknesses, and suggestions (Q28, Q29). This sequence could potentially induce a response bias, where participants’ views on the agent’s performance might have disproportionately influenced their assessment of its strengths and weaknesses. While we aimed for a comprehensive understanding of user experiences and randomized other aspects of the study including Q27 which explicitly asks for the benefits and drawbacks of each presentation method immediately after interacting with that method, future studies may benefit from randomizing the order of these questions to mitigate any potential bias.
The telemetry data extracted from Viva Insights provided valuable information on participant work patterns but did not capture a complete picture. While we focused on macro-level telemetry data such as meeting patterns, email habits, and focus time blocks, we acknowledge that micro-level data (such as moment-to-moment task switches, detailed distraction patterns, or granular activity logs) could potentially enable more fine-grained personalization. Our choice of macro-level telemetry was driven by considerations of privacy, data availability, and practical applicability in enterprise settings. However, this higher-level view might miss productivity patterns that could be captured through more detailed activity tracking. Future research could explore how combining both macro and micro-level telemetry data might enable more sophisticated personalization while carefully balancing the trade-offs between insight granularity and privacy concerns. Furthermore, there might have been external factors or variables, such as changes in workload, organizational structure, or personal circumstances, that may affect productivity and thus are essential to be considered. In addition, our system was designed to extract specific, predefined types of information from telemetry data (such as meeting patterns and email habits), with these categories hardcoded into the prompts. We did this to ensure consistent experimental conditions across participants and facilitate meaningful comparisons between different presentation methods. Since the study primarily aimed to demonstrate the feasibility and effectiveness of integrating GPT-4 with productivity tools, hence this structured approach to data extraction. While this design choice limited the system’s flexibility in discovering additional productivity patterns, we believe that with the advancing capabilities of LLMs and robust error handling, future implementations could adopt a more dynamic approach to telemetry data extraction and insight generation.
@@
In light of these limitations, it is essential to approach the study’s findings with caution and acknowledge that the results may not be universally applicable. Future research could address these limitations by employing larger and more diverse samples, extending the study duration, incorporating additional measures of productivity, and controlling for potential confounding factors. Finally, it is important to consider the extent to which the AI productivity agent’s design was influenced by the user preferences identified in Phase 1. While there is a possibility that a closer alignment might have affected the comparative results, our design was a conscientious effort to balance user needs with technical feasibility. Future research could explore a more in-depth alignment of these insights, potentially offering a refined understanding of how well the AI agent meets user expectations. This ongoing process highlights the iterative nature of developing user-centric AI tools.
@@ 8. Ethical Considerations
We adhered to strict ethical standards throughout our study to protect participant rights and privacy. All participants provided informed consent after receiving a comprehensive overview of the study’s objectives, procedures, potential benefits, and risks. Anonymity and confidentiality were maintained by using unique identification codes for participant data, securely storing all information on encrypted servers, and limiting access to authorized research team members. Our data security approach extends beyond the study period through strict data minimization practices. Following study completion, all participant contact information used for payment and study coordination was permanently deleted from our systems. The remaining anonymized research data is stored using enterprise-grade encryption and follows the organization’s established data retention policies, which include regular security assessments and updates to protect against emerging vulnerabilities. This approach ensures participant information remains protected long-term.
The use of workplace telemetry data, particularly when combined with generative AI technologies, raises significant concerns around user privacy, autonomy, and workplace surveillance  (Das Swain and Saha, 2024; Hernandez et al., 2021; Das Swain et al., 2023a). These concerns are especially pertinent as organizations increasingly adopt sophisticated tracking and analytics tools that could potentially enable intrusive monitoring practices. During the telemetry data collection process, we implemented several specific privacy-preserving measures. Our scraper tool’s interface provided participants complete visibility into the data being collected, and importantly, participants could review and modify their telemetry data before submission. We adhered to the principle of data minimization by collecting only pre-defined, high-level metrics without accessing detailed communication content, and limiting historical data to the previous month to minimize data exposure. All collected data was immediately encrypted and stored in a secure environment. This approach to telemetry data personalization was validated by our user survey findings, where 73.2% of participants expressed comfort in sharing high-level telemetry data for enhancing productivity guidance.
Beyond data collection considerations, the design of AI-based productivity systems presents additional ethical challenges. We deliberately chose to work with high-level aggregated telemetry data rather than granular activity tracking, striking a balance between providing meaningful insights and respecting user privacy. Our implementation ensures that the system serves as a supportive tool for individual productivity improvement rather than a monitoring mechanism for management oversight. The use of GPT-4 in processing workplace telemetry data introduces additional ethical considerations specific to our context. We addressed these through several design choices: utilizing the enterprise version of GPT-4 purely as an inference engine without data retention, processing only aggregated productivity metrics rather than detailed personal communications or schedules, and ensuring that all recommendations are suggestive rather than prescriptive. This approach maintains user autonomy while providing productivity support. Furthermore, we considered the impact of AI-driven productivity recommendations on work-life balance and employee well-being. While our system makes suggestions based on telemetry data, we designed it to avoid inadvertently promoting overwork or unhealthy work patterns. These ethical considerations informed both our technical implementation and evaluation approach, where we prioritized understanding user preferences and comfort levels with different types of productivity guidance. Our goal was to develop a system that enhances workplace productivity while respecting user privacy, autonomy, and well-being.
@@ 9. Conclusion
This paper presents a comprehensive study focused on understanding user preferences for productivity agents and designing personalized solutions to cater to these preferences. By conducting a user survey and analyzing the findings, we were able to identify key features and behaviors that contribute to user satisfaction and comfort when interacting with a productivity agent. Subsequently, we developed an AI-based productivity agent and compared its performance with alternative modes of telemetry data presentation, such as dashboard visualization and personalized narratives. The results from this study provide useful insights into the factors that enhance productivity and user experience, as well as areas for improvement in the design and development of personalized productivity agents. By incorporating the lessons learned from this study, future work can focus on refining and optimizing productivity-enhancing tools and solutions, ultimately leading to improved efficiency and user-centric experiences for information workers.
@@ Appendix C Prompts for Extracting Data
User Prompt for Page 1: Extract the organized meetings, invited meetings and meeting habits (along with %) and list them. If you cannot find these information, do not mention anything about it in the response. DO NOT LIST ANY NAMES OF INDIVIDUALS. List the information in this format: ‘Meetings organized in the past four weeks: X, Meetings invited to in the past four weeks: X, Meeting habits based on meetings you attended: - Advanced notice: X%, - High attendance: X%, - No overlap with other meetings: X%, - Ended on time: X%,- Joined on time: X- During working hours: X- Meeting was $\leq$ 1 hour: X- RSVP’d to invite: X- You didn’t multitask: X- Added a teams link: X%’
@@ Appendix E Prompt for Personalized Narrative and Interactive Chat Agent
Appendix H User survey responses
@@
Making progress or completing scheduled tasks 347 (95.6%)
Solving novel, unexpected tasks 278 (76.6%)
Helping a colleague accomplish a task 262 (72.2%)
Learning new techniques or information 250 (68.9%)
Staying up-to-date with messaging and emails 153 (42.1%)
Effectively managing and participating in meetings 149 (41.0%)
Establishing new connections or expanding my professional network 111 (30.6%)
Other 14 (3.9%)
@@
Yes, definitely 142 (39.1%)
Yes, somewhat 138 (38.0%)
Neutral 57 (15.7%)
No, not really 17 (4.7%)
No, not at all 9 (2.5%)
@@
Moderate impact 117 (41.8%)
Minimal impact 73 (26.1%)
Significant impact 54 (19.3%)
Unsure 27 (9.6%)
Other 9 (3.2%)
@@
Casual and friendly 46 (16.4%)
Formal and professional 23 (8.2%)
It depends on the context 95 (33.9%)
A balance of both 99 (35.4%)
No strong preference 17 (6.1%)
@@
Professional and respectful 71 (25.4%)
Direct and straightforward 69 (24.6%)
Adaptive 55 (19.6%)
Friendly and warm 50 (17.9%)
Engaging and motivational 19 (6.8%)
Unsure 12 (4.3%)
Other 4 (1.4%)
@@
Direct and to the point 99 (35.4%)
Chatty and conversational 12 (4.3%)
A balance of both 86 (30.7%)
It depends on the context 80 (28.6%)
No strong preference 3 (1.1%)
@@
Moderate Personalization 109 (38.9%)
High Personalization 100 (35.7%)
Minimal Personalization 34 (12.1%)
No personalization 15 (5.4%)
Unsure 22 (7.9%)
@@
Yes, completely comfortable 46 (16.4%)
Somewhat comfortable 106 (37.9%)
Neutral 61 (21.8%)
Uncomfortable 58 (20.7%)
Unsure 9 (3.2%)
@@
Yes, completely comfortable 121 (43.2%)
Somewhat comfortable 84 (30.0%)
Neutral 36 (12.9%)
Uncomfortable 31 (11.1%)
Unsure 8 (2.9%)
@@
Proactive suggestions 169 (60.4%)
Only specific requests 77 (27.5%)
Not at all 7 (2.5%)
Unsure 19 (6.8%)
Other 8 (2.9%)
@@
Knowledgeable 249 (88.9%)
Trustworthy 210 (75.0%)
Transparent 191 (68.2%)
Responsive 184 (65.7%)
Adaptable 172 (61.4%)
Professional 131 (46.8%)
Friendly 104 (37.1%)
Humor 92 (32.9%)
Empathetic 81 (28.9%)
No strong preference 4 (1.4%)
Other 13 (4.6%)
@@
Assistant 214 (76.4%)
Expert 157 (56.1%)
Coworker 96 (34.3%)
Mentor 93 (33.2%)
Friend 65 (23.2%)
Supervisor 13 (4.6%)
Unsure 7 (2.5%)
Other 2 (0.7%)